# On planar sections of the dodecahedron

### v3

@@ Abstract.
In the analysis of three-dimensional biological microstructures such as organoids, microscopy frequently yields two-dimensional optical sections without access to their orientation. Motivated by the question of whether such random planar sections determine the underlying three-dimensional structure, we investigate a discrete analogue in which the ambient structure is the vertex set of a Platonic solid and the observed data are congruence classes of planar intersections. For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq\Pi\cap V\subseteq V$, and ask whether this statistic determines $X$ up to isometry. We show that this is not the case: there exist two non-isometric $7$-element subsets with identical planar statistics.
@@ 1. Introduction
High-resolution imaging techniques in biological systems frequently produce planar sections of inherently three-dimensional structures. A notable example appears in organoid research, where multicellular aggregates are cultivated in three-dimensional media, yet standard microscopy visualizes them in two-dimensional optical slices, [2, 3, 6]. When we think of one two-dimensional slice through a three-dimensional space filled with more or less identical objects each in a random position and orientation, then we effectively obtain a distribution of random slices of one fixed object without information on what slice exactly we are looking at. More precisely, the resulting observed data consists of a distribution of planar point configurations or shapes up to planar congruence. A fundamental question, both theoretically and for interpretation of imaging data, is whether this distribution encodes the underlying three-dimensional structure uniquely. The problem is rather different from problems in $X$-ray image analysis or geometric tomography, where the position of the intersecting plane is part of the data that can be analysed.
@@
The aim of this work is to study this question in a highly controlled discrete setting. We fix a finite vertex set $V$ of a Platonic solid in $\mathbb{R}^{3}$ and consider subsets $X\subset V$ as discrete models for three-dimensional structures. Planes intersecting $V$ produce subset inclusions $\Pi\cap X\subseteq V$ of planar sets. Only the Euclidean congruence types of these planar configurations are assumed observable, and each plane is unlabelled. Thus the observable information associated to $X$ is merely the distribution of planar congruence classes of all such intersections.
We refer to this distribution as the planar statistic of $X$, see Definition 1. Our primary question is whether two subsets with identical planar statistics must be isometric in $\mathbb{R}^{3}$. By analogy to results of Mallows and Clark [4] on intercept distributions in the plane, non-uniqueness is expected. We establish such non-uniqueness in the case of the regular dodecahedron, see Theorem 3.
This also implies that there exist two three-dimensional convex bodies whose distribution of isometry classes of two-dimensional intersections is identical, while the two convex bodies are not themselves isometric, see Theorem 4.
@@
The geometric step then promotes this combinatorial configuration to a pair of convex subsets of the plane. To each side in $X$, respectively in $Y$, one attaches a small congruent isosceles triangular hat pointing outward, thereby turning the octagon into two convex dodecagons $K_{X}$ and $K_{Y}$. The caps are chosen so small that every chord of $K_{X}$ or $K_{Y}$ either coincides with a chord of the original octagon, or differs from it only in a tiny segment inside one of the caps. The contribution of a cap to the chord–length distribution depends only on which side it sits on, so the matching of unordered pairs of sides in $X$ and $Y$ forces the chord–length distributions of $K_{X}$ and $K_{Y}$ to coincide, although the polygons themselves are not congruent.
The present dodecahedral construction follows the same two-step pattern in a three-dimensional setting. Here the regular octagon is replaced by the regular dodecahedron $D$ with vertex set $V$, and the role of chords is taken over by planar sections through vertex–planes. The first, combinatorial step is the search for two subsets $S,T\subset V$ with identical planar statistics in the sense of Definition 1. Equivalently, for every vertex–plane $\Pi$ the inclusion $\Pi\cap S\subset V$ has, up to isometry of $D$, the same type and multiplicity distribution as $\Pi\cap T\subset V$. Theorem 3 shows that such a non-congruent pair $(S,T)$ exists inside the vertex–plane incidence structure of the dodecahedron.
In the second step we again encode this discrete data into convex bodies by a local modification at selected vertices. Starting from $D$, we cut off a small congruent triangular pyramid at each vertex $v\in X$, obtaining a polytope $K_{X}$ for $X\subseteq V$. When an affine plane $H$ meets $D$ near some vertices, the local shape of the section $K_{X}\cap H$ in a neighbourhood of those vertices is completely determined, up to isometry, by the configuration $\Pi\cap X$, where $\Pi$ is the nearby vertex–plane through the intersected vertices. The distribution of isometry classes of planar sections of $K_{X}$ can therefore be written as a convex combination of contributions coming from the finitely many planar intersection types $\Pi\cap X\subset\Pi\cap V\subseteq V$, with coefficients prescribed precisely by the planar statistic $\mathrm{PS}(X)$.
From this viewpoint, Theorem 4 is the exact three-dimensional analogue of the Mallows–Clark theorem: the equality $\mathrm{PS}(S)=\mathrm{PS}(T)$ ensures that the section distributions of $K_{S}$ and $K_{T}$ coincide, while the underlying subsets $S$ and $T$ and hence the patterns of truncated vertices of $D$ are not related by any isometry.
Beyond the discrete-geometric questions addressed here, our section statistics interface naturally with classical inverse problems in geometric tomography and integral geometry [1, 7], where distributions of slices, chords, and projections play a central role. Related stereological ideas go back at least to Wicksell’s corpuscle problem [8] and remain fundamental in modern microscopy when only section data are available.
@@ 3. Convex bodies
where $\mathcal{A}(3,2)$ is the affine Grassmannian of planes in $\mathbb{R}^{3}$. Note that $\mathcal{A}(3,2)$ carries a standard infinite rotation- and translation-invariant measure $\lambda$ and we call the Borel measure $(\Phi_{K})_{*}(\lambda)$ on $\mathcal{P}$ the distribution of isometry classes of planar sections. Our normalization of $\lambda$ is in such a way that $(\Phi_{K})_{*}(\lambda)(\mathcal{P}\setminus\{\varnothing\})=W_{1}(K)$ for a compact convex set $K$, where $W_{1}$ is the first quermassintegral.
@@ Acknowledgments
We thank Vadim Alekseev for interesting discussions on this topic.

### v4

@@ Abstract.
In the analysis of three-dimensional biological microstructures such as organoids, microscopy frequently yields two-dimensional optical sections without access to their orientation. Motivated by the question of whether such random planar sections determine the underlying three-dimensional structure, we investigate a discrete analogue in which the ambient structure is the vertex set of a Platonic solid and the observed data are congruence classes of planar intersections. For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq\Pi\cap V\subseteq V$, and ask whether this statistic determines $X\subset V$ up to isometry. We show that this is not the case: there exist two non-isometric $7$-element subsets with identical planar statistics.
@@ 1. Introduction
High-resolution imaging techniques in biological systems frequently produce planar sections of inherently three-dimensional structures. A prominent example appears in organoid research, where multicellular aggregates are cultivated in three-dimensional media, yet standard microscopy visualizes them in two-dimensional optical slices, [4, 5, 11]. In many experimental pipelines, one does not obtain a calibrated full-volume scan, but rather a large collection of optical sections or histological slices taken at (approximately) random offsets and orientations.
From a mathematical viewpoint, such data amount to a distribution of planar configurations, observed only up to planar congruence. This leads to a basic inverse question: does the distribution of random planar sections identify the underlying three-dimensional structure uniquely? The problem is conceptually different from classical geometric tomography, where the position of the slicing plane is part of the measured data; here we treat the slice as unlabelled, and only the isometry class of the resulting planar configuration is retained.
The classical entry point in stereology is Wicksell’s corpuscle problem for spheres [15]. For a single sphere of radius $r$, a plane at offset $u\in[-r,r]$ cuts a disk of radius
$\rho=\sqrt{r^{2}-u^{2}}\in[0,r].$
If the plane offset is sampled uniformly, the induced (non-normalized) section-radius measure on $(0,r)$ is
$\mu_{r}(d\rho)=\frac{2\rho}{\sqrt{r^{2}-\rho^{2}}}\,d\rho,\qquad\mu_{r}([0,r])=2r.$
Passing from a single sphere to a population, let $f(r)$ be the density of sphere radii in $\mathbb{R}^{3}$ and let $g(\rho)$ be the density of observed section radii. Then $g$ is related to $f$ by an Abel-type integral transform [15, 3]:
$g(\rho)=2\rho\int_{r=\rho}^{\infty}\frac{f(r)}{\sqrt{r^{2}-\rho^{2}}}\,dr,\qquad\rho>0,$
and one can invert this relation and recover $f$ from $g$ via
$f(r)=-\frac{1}{\pi}\frac{d}{dr}\int_{\rho=r}^{\infty}\frac{g(\rho)}{\sqrt{\rho^{2}-r^{2}}}\,d\rho,\qquad r>0.$
In particular, in the spherical case the radius distribution is uniquely determined by the section statistics.
Beyond spheres, identifiability becomes subtle. For particles of the form $sK$ (random orientation and scale $s$), section functionals such as area lead to one-dimensional integral equations, and the scale distribution can often be identified [12, 14]. Related stereological methods and estimators also exist for polyhedral shapes such as cubes [9, 10]. On the other hand, there are genuine non-identifiability results: Cruz-Orive, solving a conjecture of Moran, exhibited distinct distributions of ellipsoids with identical induced distributions of unoriented planar sections [1, 8].
@@
The aim of this work is to study such identifiability questions at first in a highly controlled model, motivated by random-slice data in microscopy. We fix a finite vertex set $V$ of a Platonic solid in $\mathbb{R}^{3}$ and regard subsets $X\subset V$ as discrete proxies for three-dimensional microstructures. Each vertex-plane $\Pi$ determines an inclusion $\Pi\cap X\subseteq\Pi\cap V\subseteq V$, and the only observable information is the Euclidean congruence type of this inclusion. Aggregating over all vertex-planes, we obtain a distribution of planar congruence classes; we call it the planar statistic of $X$, see Definition 1.
Our guiding question is whether the planar statistic determines $X$ up to isometry in $\mathbb{R}^{3}$. In analogy with non-uniqueness phenomena for intercept and chord-length distributions in the plane [6], one expects that low-dimensional section statistics can fail to be injective. Our main theorem confirms this expectation in the case of the regular dodecahedron: we exhibit two non-congruent $7$-element subsets $S,T\subset V$ with identical planar statistics, see Theorem 3. In particular, even in an idealized, noise-free regime with isotropic plane sampling, uniformly sampled slices need not suffice to distinguish the underlying structure.
As a geometric consequence, we construct two polytopes in $\mathbb{R}^{3}$ whose distributions of isometry classes of unoriented planar sections coincide, although the bodies themselves are not isometric; see Theorem 4. So, while the result of Cruz-Orive [1] shows non-identifiability for distributions of convex sets, our construction yields non-identifiability already for single polytopes. This result is an analogue of classical non-uniqueness phenomena in geometric tomography [2, 13].
@@
The geometric step then promotes this combinatorial configuration to a pair of convex subsets of the plane. To each side in $X$, respectively in $Y$, one attaches a small congruent isosceles triangular hat pointing outward, thereby turning the octagon into two convex dodecagons $K_{X}$ and $K_{Y}$. The contribution of a cap to the chord–length distribution depends only on which side it sits on, so the matching of unordered pairs of sides in $X$ and $Y$ forces the chord–length distributions of $K_{X}$ and $K_{Y}$ to coincide, although the polygons themselves are not congruent.
The present dodecahedral construction follows the same two-step pattern in a three-dimensional setting. Here the regular octagon is replaced by the regular dodecahedron $D$ with vertex set $V$, and the role of chords is taken over by planar sections through vertex–planes. The first, combinatorial step is the search for two subsets $S,T\subset V$ with identical planar statistics in the sense of Definition 1. Equivalently, for every vertex–plane $\Pi$ the inclusion $\Pi\cap S\subset\subset\Pi\cap V\subset V$ has, up to isometry of $D$, the same type and multiplicity distribution as $\Pi\cap T\subset\Pi\cap V\subset V$. Theorem 3 shows that such a non-congruent pair $(S,T)$ exists inside the vertex–plane incidence structure of the dodecahedron.
In the second step we again encode this discrete data into convex bodies by a local modification at selected vertices. Starting from $D$, we cut off a small congruent triangular pyramid at each vertex $v\in X$, obtaining a polytope $K_{X}$ for $X\subseteq V$. When an affine plane $H$ meets $D$ near some vertices, the local shape of the section $K_{X}\cap H$ in a neighbourhood of those vertices is completely determined, up to isometry, by the configuration $\Pi\cap X\subset\Pi\cap V\subset V$ , where $\Pi$ is the nearby vertex–plane through the intersected vertices. The distribution of isometry classes of planar sections of $K_{X}$ can therefore be written as a convex combination of contributions coming from the finitely many planar intersection types $\Pi\cap X\subset\Pi\cap V\subseteq V$, with coefficients prescribed precisely by the planar statistic $\mathrm{PS}(X)$.
From this viewpoint, Theorem 4 is (upon replacing faces by vertices)the exact three-dimensional analogue of the Mallows–Clark theorem: the equality $\mathrm{PS}(S)=\mathrm{PS}(T)$ ensures that the section distributions of $K_{S}$ and $K_{T}$ coincide, while the underlying subsets $S$ and $T$ and hence the patterns of truncated vertices of $D$ are not related by any isometry.
@@ 3. Convex bodies
where $\mathcal{A}(3,2)$ is the affine Grassmannian of planes in $\mathbb{R}^{3}$. Note that $\mathcal{A}(3,2)$ is naturally a fibre bundle over $\mathbb{R}{\rm P}(2)$ with fibre $\mathbb{R}$, where the fibre over a line $u\in\mathbb{R}{\rm P}(2)$ is the set of all planes with normal $u$, where the element in the fibre determines the offset. A natural measure $\lambda$, invariant under isometries of $\mathbb{R}^{3}$, can then be constructed as the product of the surface area measure on $\mathbb{R}{\rm P}(2)$ and the Lebesgue measure on the fibre $\mathbb{R}$. With this normalization of $\lambda$, we obtain $(\Phi_{K})_{*}(\lambda)(\mathcal{P}\setminus\{\varnothing\})=6\cdot W_{2}(K)$ for a compact convex set $K$, where $W_{2}$ is the second quermassintegral. We call the Borel measure $(\Phi_{K})_{*}(\lambda)$ on $\mathcal{P}$ the distribution of isometry classes of planar sections.
@@ Acknowledgments
We thank Vadim Alekseev and Leon Renkin for interesting discussions on this topic.